OpenStudy (anonymous):

Factor the polynomial completely: 3a^4 - 8a^3 + a^2

OpenStudy (anonymous):

Start by pulling out as many a's as possible: a^2(3a^2-8a+1) Then look at the inside and notice it is a quadratic equation, which you most likely have factored by itself before: 3a^2-8a+1 = (3a +- ?)(a +- ?) [I was able to assume that the coefficients of the a's are 3 and 1 because that is the only possible combination that multiplies to 3a^2. This little time-saver works whenever the coefficient of a^2 is a prime number, in which case one of the a's will be the coefficient itself while the other is just 1]

OpenStudy (aum):

3a^4 - 8a^3 + a^2 = a^2(3a^2-8a+1) That is it. The quadratic part does not factor nicely and so you leave it as it is.

OpenStudy (anonymous):

I just assumed since it said "factor completely" that it wanted us to use the quadratic formula if necessary

OpenStudy (anonymous):

Thanks. This is getting me confused because im always told to first find the GCF, there is none so I'm always stuck after that.

OpenStudy (aum):

Yes, find GCF first: 3a^4 - 8a^3 + a^2 every term has a^2 as a factor and it is the GCF that can be factored out: a^2(3a^2 - 8a + 1). Then see if the quadratic part can be factored. In this case it does not factor and so leave it as it is.